OpenStudy (anonymous):

lim t->0 from the right of (1-Ln(t))/e^(1/t)

OpenStudy (anonymous):

apply l'hospital rule and simplify.

OpenStudy (anonymous):

0

OpenStudy (anonymous):

i did, got 1/t// e^1/t*(-t^-2)... the series gets larger and larger

OpenStudy (anonymous):

from the left = infinity

OpenStudy (anonymous):

Cyter why do you say 0?

OpenStudy (angela210793):

unless i'm wrong the lim is infinite/infinite so as Brackett said we use Hospital's rule lim x--> 0 from the right of [(1/t)/(-e^1/t^1/t^2)=0

OpenStudy (anonymous):

the series got bigger and bigger because u applied l'hospital wrongly...u need to stop applyin l'hospital, wen u don't hav its 0/0 or infinity/infinity form...u'll get 0.

OpenStudy (anonymous):

i cant recall, is 0/inf a legitimate answer? don't I need L'H again?

OpenStudy (anonymous):

you need l'hopital's rule when you only have an indeterminate form as the limit

OpenStudy (anonymous):

0/infinity = 0

OpenStudy (anonymous):

i thought 0/inf was indeterminate....I think thats my problem

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

see wenever u come across any indeterminate form, think for urself , wy can't u determine its value. take the case of 0/inf we hav 0/2=0/17=0/10239870870 as 0. then wy should u hav doubt on 0/inf. dividing reduces value, wen u divide by inf. it totally reduces the value to 0. don't get confused by the presence of o and inf together :)

OpenStudy (anonymous):

Excellent explanation Brackett, i was indeed getting confused.